OpenStudy (anonymous):

Blake has entered a mountain bike competition at Starvation Flats Bike Park. Placement in the competition is determined by a person's average speed over four laps around the course. The course itself is 2/5 of a mile long. If Blake completed the four laps in 12 minutes, then his average speed around the track is miles per hour.

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (texaschic101):

speed = distance / time speed = (2/5) / 12 speed = 2/5 * 1/12 speed = 2/60 = 1/30 um...I think it is 1/30 of a mile per hr.

OpenStudy (anonymous):

No. It was 8 MPH

OpenStudy (funnyguy2):

2/5 * 1/12 = 0.2

OpenStudy (anonymous):

Each lap around the track is of a mile long, so four laps around the track is as follows. Blake completed the four laps in 12 minutes, which needs to be converted to hours. To find Blake's speed in terms of miles per hour divide the distance miles by of an hour. Therefore, Blake's average speed around the track is 8 miles per hour.